1 Which key principal explains the positively sloped portion of the short-run marginal cost curve? 2....

1

Which key principal explains the positively sloped portion of the short-run

marginal cost curve?

2.

What is happening to marginal cost as marginal product is rising?

3.

Comment on the following statement:

Assume a firm’s short-run marginal cost is less than the short-run average cost. If the

firm increases its output level, will the firm’s average cost increase or decrease?

Explain relative to marginal cost.
Answer #1

Answer

(1)

The key principle that explains the positively sloped portion of the short-run marginal cost curve is Law of Variable proportion.

In short run one factor is fixed(like capital is fixed), According to law of variable proportion when we continue hiring variable factor like labor keeping Capital fixed there will definitely come a point when Marginal productivity of labor starts decreasing i.e. addition to output when we hire extra unit of labor will decline.

(2)

Suppose marginal product is rising. This means that extra output produced when we hire one additional unit of labor will increase. As Marginal cost is the additional cost required to produce one additional unit of output and as discussed above that one additional unit of labor will produce more output because of increasing marginal product. Thus In order to produce 1 additional unit of output lesser labor will required and hence marginal cost will be less.

Thus as marginal product is rising, Marginal cost will be falling.

(3)

AC = TC/Q , where AC = Average cost , TC = Total cost and Q = quantity

AC will be falling if d(AC)/dQ < 0 and AC will be rising if d(AC)/dQ > 0

d(AC)/dQ = d(TC/Q)/dQ = (Q(d(TC)/dQ) - TC)/Q2

= (d(TC)/dQ)/Q - (TC/Q)/Q

= MC/Q - AC/Q

= (MC - AC)/Q

Here MC < AC

=> d(AC)/dQ = (MC - AC)/Q < 0

=> d(AC)/dQ < 0 => AC will be falling.

Hence,

If the Firm increases its output level, the firm’s average cost will decrease